# AI and the Net-Zero Journey: Energy Demand, Emissions, and the Potential for Transition

### v1

@@ Abstract
The emergence of new technologies inevitably brings inherent challenges and a degree of skepticism regarding claimed benefits and practical applications. This is a natural part of technological evolution. In this article, we focus mainly on the impact of growing infrastructure needs for AI, specifically data centers, and the increasing energy demand required to power them. While the energy mix that will supply these data centers is important and is expected to evolve in the coming years, our main emphasis is on CO2 emissions and their broader climate impact. We will especially explore the interplay between the rise of AI, data centers, energy demand, and emissions. In determining whether AI results in a net positive or negative contribution to emissions, it’s essential to consider scenarios both with and without dedicated emissions reduction measures. Additionally, we consider the carbon benefits directly attributed to AI across various sectors of energy industry. Quantifying these elements presents significant challenges and uncertainties, given the wide variation in estimates and forecasts from both public and private institutions. Nevertheless, we have made reasonable assumptions and considered existing paradigms, such as the potential for AI to reduce an organization’s GHG emissions or achieve global CO2 reductions presented in [Maher et al. (2022)]. Our analysis mainly draws on the energy scenarios presented in [She (2025)], which outlines global energy mix trends and the anticipated impact of technology on reducing emissions.
@@ 1 Introduction
The concept of net-zero greenhouse gas (GHG) emissions is crucial for combating climate change and limiting global warming to well below 2°C, ideally to 1.5°C pre-industrial levels, as outlined in the Paris Agreement [Nations (2015)]. Achieving net-zero means balancing the total amount of GHGs released into the atmosphere with an equivalent amount being removed, thus stopping the accumulation of these gases that drive climate change. This goal requires significant transformations across all sectors of the economy, including energy, transportation, manufacturing, and agriculture [She (2025), Benes et al. (2024)]. This necessitates a comprehensive strategy and a multifaceted approach involving significant change in energy systems, agriculture and land use, and in industrial processes. The sheer scale of this economic transformation is a major challenge [Krishnan et al. (2022)]. For example, to reach net zero by 2050, Shell scenarios suggest that \$3-4 trillion of commercially viable investment in low-carbon energy is required each year up from current \$1.7 trillion a year investment [She (2024)]. While these are large numbers, the difference from a reasonable baseline energy system investment profile is much smaller.
@@
Generating electricity is a growing contributor to climate change, primarily because coal, oil, and gas are and will still be widely used compared to cleaner energy sources [She (2025)]. While electricity demand has been steady over the years, it has recently surged and is projected to continue increasing its share until 2050. This growth is driven by the expanding digital economy and the rise of artificial intelligence (AI). Projections indicate a substantial surge in power consumption from data centers, which are essential for AI infrastructure. Some forecasts suggest that electricity demand from this sector could double by 2026 [Report (2024a)]. The increasing complexity and computational intensity of AI models, especially generative AI, require significantly more energy per query compared to traditional internet services. For instance, a single ChatGPT query consumes approximately ten times the electricity of a typical Google search [de Vries (2023)]. Although the numbers these estimates are based have already been outdated, and the AI assistants can generate more information than just word tokens, it is extensively quoted and generally accepted that queries to AI models consume more energy than simple Google searches. This rising demand puts considerable pressure on existing electricity infrastructure, necessitating a re-evaluation of energy production and consumption patterns [Norris et al. (2025); Report (2024b)].
@@
Major tech companies and cloud providers like Microsoft and Google have reported that their emissions have increased recently [goo (2024); mic (2024)]. This trend is expected to continue due to the rush towards building more data centers in the next few years to support the rising AI demand and the ambition to become dominant players in the AI. Several reports, including the IEA report on AI and power demand trends, indicate that GHG emissions will inadvertently rise in the near future as the alternative energy sources are expected to take more time to become significant contributors [She (2025)].
@@
However, AI also presents a transformative opportunity for accelerating the transition to a net-zero future. AI’s unique ability to process vast amounts of complex data on energy systems, emissions, and climate impact allows for more informed and data-driven approaches to decarbonization. AI can optimize energy grids, improve the efficiency of renewable energy sources through better forecasting, and support the development of more sustainable industrial processes and materials. Additionally, AI can enhance energy efficiency, availability and reduce waste across various sectors, from transportation and buildings to industrial manufacturing, significantly contributing to emissions reduction. Some reports suggest that AI alone could potentially reduce global GHG emissions by 4% to 16% by 2030 [She (2025); Maher et al. (2022)], and 1.4 gigatonnes by 2035 in the widespread adoption case [2024 (2025)].
@@
The advancements in AI for science (see, e.g., [Chandra et al. (2024)]) demonstrate that AI can significantly accelerate complex challenges such as weather modeling and prediction, climate modeling, and CCS storage modeling technologies, potentially by up to 100,000 times. At Shell, we have conducted projects investigating the role of AI in CCS within both traditional aquifers and land use. Our findings indicate that AI can greatly expedite the process of identifying suitable sites and accelerating project execution by several orders of magnitude. We anticipate that in the next 5-10 years, these areas will present significant potential and opportunities to offset the carbon emissions [Pawar et al. (2025)].
@@ 2 Methodology
By employing MCMC simulations with these probabilistically defined inputs, we can generate a range of potential emission scenarios for both the near and long term. This approach allows for a more nuanced understanding of the uncertainties involved and enables sensitivity analysis based on the factors with the most significant influence on future data center and AI-related emissions. The ultimate goal is to provide robust, probabilistic insights that can inform strategies for mitigating the environmental impact of this rapidly growing sector.
@@ 3 The AI Boom and its Energy Appetite
Figure 1 shows the number of foundational models with more than 10 billion parameters built between 2021 and 2024 [data courtesy eco (2025)]. It also illustrates the increase in revenues from high-density GPUs from 2022 and the projection for 2028 (Source: ele (2024)). As can be seen, more investments are being made towards building GPU servers to support the development of large AI models. Although NVIDIA GPUs dominate the market for chips and low-level software libraries, there are no barriers for new entrants to join the arena of building new foundational models. Due to supply and demand mismatches and other approval process-related issues, commissioning grid infrastructure has temporarily slowed the pace of data center growth. However, the overall trend remains one of rapid expansion.
Figure 1: The figure illustrates the growth of foundational models between 2021 and 2024, and the revenue growth of high-density GPU markets. Data courtesy from eco (2025)
@@
However, this progress comes with significant consequences: a rapidly increasing demand for computational resources required to support the tasks in the life cycle of AI models, which involves both training (a one-time or occasional effort) and inference (repeatedly executed when the model is deployed for consumption). Although the work done (measured in FLOPs) per dollar and per watt has increased, meaning the power consumption per FLOP has decreased (see Figure 2 (left)), the number of FLOPS required to train AI models has continued to rise (see Figure 2 (right)). Consequently, the total power consumption for training AI models keeps increasing. This translates directly into escalating energy consumption [Report (2025)]. For instance, it was reported in [Patterson et al. (2021)] that training the large language model (LLM) powering ChatGPT-3 consumed almost 1300 megawatt hours of energy, equivalent to the annual energy usage of 130 American homes. Furthermore, an analysis by OpenAI suggests that the amount of power needed to train AI models has grown exponentially since 2012, doubling roughly every 3.4 months as the models become more sophisticated [ope (2018)].
@@
Computational Demands of Training and Inference: AI models, especially deep learning models and LLMs, require immense computational power for both training and inference [de Vries (2023)].
@@
  • Inference is the deployment and utilization of the trained AI models in real-world applications, where the model interprets new data and generates outcomes or predictions. Despite innovations in model efficiency, the inference phase accounts for the majority (60%) of the energy footprint [Report (2024b)] (see Table 1 for electricity consumption of various tasks that AI chatbots are typically asked to execute).
@@ 4 Data Centers: The Epicenter of AI’s Energy Demand
Data centers are the backbone of the AI ecosystem, serving as the primary energy consumers that power the computational workloads required for training and deploying AI models. These facilities house vast arrays of servers, storage devices, and networking equipment, all of which demand significant amounts of electricity to operate. The rapid expansion of AI, particularly deep learning and large language models, has led to an unprecedented surge in data center growth, intensifying concerns about their environmental impact [Report (2024b)].
The data center industry is experiencing rapid growth worldwide, driven by the increasing demand for cloud computing, digital services, cryptocurrencies, and AI applications. The number of data centers is increasing, and their sizes are also expanding. As of March 2024, there were approximately 9580 data centers globally, with more than half of them located in the United States [dat (2025)]. Between 2015 and 2024, the electricity consumption at data centers increased from under 200 TWh to over 400 TWh [2024 (2025)]. Data centers are projected to consume between 4.6% and 9.1% of U.S. electricity generation annually by 2030, up from an estimated 4% today. Some forecasts even predicted that data centers could draw up to 21% of the world’s electricity supply by 2030 [Jones (2018)]. In the US, fifteen states account for 80 percent of the national data center load, with data centers comprising a quarter of Virginia’s electric load in 2023. New data centers are being built with capacities ranging from 100 to 1000 megawatts, equivalent to the load from 80,000 to 800,000 homes.
In [Report (2024b)], the authors have developed four growth scenarios for US data center electricity load from 2023 to 2030: low growth (3.7 percent annual), moderate growth (5 percent annual), high growth (10 percent annual), and higher growth (15 percent annual). In the higher growth scenario, data center loads are projected to increase by a staggering 166% from 152 TWh/year in 2023 to 403 TWh/year in 2030, with the majority of this demand attributed to AI. Such projections highlight the significant energy demands of AI and data centers and the need for innovative energy management and infrastructure planning to accommodate this growth.
@@
Despite the growing emphasis on renewable energy, many regions still rely heavily on fossil fuels, especially, natural gas, to power their electricity grids [Rotman (2025)]. This may mean that a significant portion of data centers’ energy consumption is directly contributing to GHG emissions. In regions with limited access to clean energy sources, the increasing electricity demands of data centers can exacerbate the reliance on fossil fuels, hindering efforts to decarbonize the energy sector. For instance, in [Rotman (2025)], the author quoted that natural gas is the only affordable choice given the need to meet the 24-7 electricity demand from the huge data center.
@@
As AI has become increasingly pervasive, it is difficult to estimate how much electricity consumption can be attributed solely to AI-related workloads. Some studies using proxies estimate that AI accounts for 10% to 20% of total data center electricity consumption [see 2024 (2025) for example], but this percentage is rapidly increasing. AI workloads, including computation, storage, and networking, have specific energy needs within data centers [Report (2024b, a)].
@@ 5 The Energy Consumption Scenarios and Near-Term Outlook for Data Centers and AI
Shell’s 2025 Energy Security Scenarios explore AI’s transformative role across different pathways, each highlighting distinct impact on energy systems and societal change [She (2025)]. In the Surge scenario, AI is a key driver of economic growth and societal transformation. AI optimizes manufacturing processes, enhances productivity across various sectors, and leads to increased economic activity. This surge in economic activity results in higher energy demand, which is managed efficiently by AI-powered smart grids. Additionally, AI accelerates the transition to low-carbon energy sources by optimizing the integration of renewable energy and enhancing energy storage solutions. The Archipelagos scenario presents a different picture, where AI development is hindered by global concerns about resource security, border control, and trade restrictions. Geopolitical tensions limit the sharing of AI technologies and data, slowing progress. As a result, technological advancements are slower, impacting the transformation of energy systems. In the Horizon scenario, AI plays a crucial role in achieving net-zero emissions by 2050. AI-driven innovations in energy efficiency and carbon capture technologies are pivotal in reaching this goal. AI also helps limit global warming to 1.5°C by optimizing energy consumption and reducing GHG emissions (Figure 4).
Figure 4: Electricity generation and associated CO2 emissions between 2020-2060 under the three Shell Energy Security scenarios. Of those, only Horizon is reaching Net Zero by the 2050s. [Courtesy of She (2025)].
@@
Building on the insights from Shell’s Energy Securities scenarios, the Energy and AI 2025 report by the International Energy Agency (IEA) [2024 (2025)] outlines four sensitivity cases to explore different scenarios for AI uptake and its impact on the energy sector (Figure 5a). The Lift-Off Case assumes higher rates of AI adoption and proactive measures to reduce energy sector bottlenecks, envisioning rapid expansion of AI technologies and infrastructure. This scenario highlights the transformative potential of AI when supported by robust energy policies and investments, leading to significant advancements in AI capabilities and widespread adoption. The Headwinds Case incorporates various bottlenecks, including macroeconomic challenges and delays in building out necessary energy infrastructure, representing a more cautious and constrained growth scenario. This case underscores the importance of addressing external factors that can hinder AI development, such as economic instability and infrastructure limitations, which can slow progress and reduce the overall impact of AI.
@@
Figure 5: a) Four scenarios for electricity demand by data centers globally between 2020-2035 (reproduced from 2024 (2025)); b) Plausible scenarios for electricity usage by AI applications between 2020 and 2035. (reproduced from Paccou and Wijnhoven (2024)).
@@
In a recent study [Paccou and Wijnhoven (2024)], the authors presented a comprehensive outlook using a system dynamics approach to model four scenarios for AI electricity consumption: Sustainable AI, Limits To Growth, Abundance Without Boundaries, and Energy Crisis. These scenarios provide insights into potential future trajectories of AI electricity use, influenced by current AI infrastructure decisions. While all scenarios show initial growth, their paths diverge after 2030 due to various internal and external factors (Figure 5b). The Sustainable AI scenario prioritizes efficiency and sustainability, with energy consumption increasing from 100 TWh in 2025 to 785 TWh in 2035. This scenario balances technological advancement with environmental stewardship. In contrast, the Limits To Growth scenario shows minimal increase in consumption, from 510 TWh in 2030 to 570 TWh in 2035, indicating stunted economic expansion due to constraints like power availability and data scarcity. The Abundance With- out Boundaries scenario projects unchecked growth, with consumption reaching 1,370 TWh by 2035. This scenario highlights risks such as power centralization and inequitable AI access. Meanwhile, the Energy Crisis scenario underscores the risks of mismatched energy demand and infrastructure, with consumption peaking at 670 TWh before dropping to 190 TWh by 2035, indicating potential energy crisis.
@@
Specific insights from these scenarios include the dominance of generative AI in the Sustainable AI scenario, focusing on less energy-intensive models. The Limits To Growth scenario highlights constraints in AI development due to power, chip manufacturing, and data scarcity. The Abundance Without Boundaries scenario warns of risks associated with unchecked AI growth, including unsustainable infrastructure costs, global inequality, and increased e-waste. The Energy Crisis scenario emphasizes the likelihood of underestimated future electricity needs due to insufficient grid planning and inaccurate demand projections. The authors note that projections for electricity consumption in TWh face challenges due to the nascent nature of the discipline, making direct comparisons difficult. However, some convergence is observed: the AI Abundance Without Boundaries scenario projects 880 TWh usage by 2030, aligning with estimates by [Patel et al. (2024)]. Another projection by [Kindig (2024)] of 652 TWh by 2030 aligns with the Sustainable AI scenario, suggesting a balanced growth trajectory.
@@
To provide a comprehensive understanding of the potential futures for AI development and energy consumption, we have matched the Shell scenarios with the IEA sensitivity cases and the AI scenarios from Paccou and Wijnhoven (2024). This alignment is based on the shared themes and underlying assumptions of each scenario. The Shell scenarios offer a broad perspective on global economic and technological trends, while the IEA sensitivity cases focus on the specific impacts of AI on the energy sector. The AI scenarios from Paccou and Wijnhoven (2024) provide detailed insights into the potential trajectories of AI electricity consumption. By matching these scenarios, we can explore how different factors such as geopolitical tensions, technological advancements, and sustainability efforts might influence the future of AI and energy. This integrated approach allows us to identify common challenges and opportunities, offering a more nuanced and holistic view of the interplay between AI development and energy consumption:
@@
    Surge - Lift-Off Case - Abundance Without Boundaries: The Surge scenario envisions rapid AI adoption and technological advancements, driving significant economic growth. Similarly, the Lift-Off Case from the IEA assumes high AI adoption rates and proactive measures to reduce energy bottlenecks, highlighting AI’s transformative potential. The Abundance Without Boundaries scenario by Paccou and Wijnhoven (2024) projects unrestricted AI growth, emphasizing its power to drive economic and technological progress without major constraints. These scenarios share a vision of rapid and expansive AI growth fueling economic and technological advancements. Under these scenarios, the energy demand by data centers continues to grow past 2035 exceeding 1700 TWh. AI accounts for more than three quarters of the electricity usage (Figure 6a).
@@
    Archipelagos - Headwinds Case - Limits To Growth: The Archipelagos scenario depicts technology development hindered by global concerns about resource, border, and trade security, leading to fragmented efforts and slower progress. Similarly, the Headwinds Case from the IEA includes various bottlenecks, such as macroeconomic challenges and infrastructure delays, representing cautious growth with significant obstacles. The Limits To Growth scenario by Paccou and Wijnhoven (2024) highlights constraints on AI development due to resource limitations, emphasizing the need for careful management of AI’s energy demands. These scenarios share a focus on limitations and challenges that hinder AI and energy sector progress. Under these scenarios, energy demand by data centers plateaus by 2030 at around 700 TWh. AI accounts for approximately 80% of the electricity usage (Figure 6b).

### v2

@@ Abstract
The emergence of new technologies inevitably brings inherent challenges and a degree of skepticism regarding claimed benefits and practical applications. This is a natural part of technological evolution. In this article, we focus mainly on the impact of growing infrastructure needs for AI, specifically data centers, and the increasing energy demand required to power them. While the energy mix that will supply these data centers is important and is expected to evolve in the coming years, our main emphasis is on CO2 emissions and their broader climate impact. We will especially explore the interplay between the rise of AI, data centers, energy demand, and emissions. In determining whether AI results in a net positive or negative contribution to emissions, it’s essential to consider scenarios both with and without dedicated emissions reduction measures. Additionally, we consider the carbon benefits directly attributed to AI across various sectors of energy industry. Quantifying these elements presents significant challenges and uncertainties, given the wide variation in estimates and forecasts from both public and private institutions. Nevertheless, we have made reasonable assumptions and considered existing paradigms, such as the potential for AI to reduce an organization’s GHG emissions or achieve global CO2 reductions presented in [bcg_climate_2022]. Our analysis mainly draws on the energy scenarios presented in [Shell_Sce_2025], which outlines global energy mix trends and the anticipated impact of technology on reducing emissions.
@@ 1 Introduction
The concept of net-zero greenhouse gas (GHG) emissions is crucial for combating climate change and limiting global warming to well below 2°C, ideally to 1.5°C pre-industrial levels, as outlined in the Paris Agreement [paris_agreement]. Achieving net-zero means balancing the total amount of GHGs released into the atmosphere with an equivalent amount being removed, thus stopping the accumulation of these gases that drive climate change. This goal requires significant transformations across all sectors of the economy, including energy, transportation, manufacturing, and agriculture [Shell_Sce_2025, DOE2024_AI4Energy]. This necessitates a comprehensive strategy and a multifaceted approach involving significant change in energy systems, agriculture and land use, and in industrial processes. The sheer scale of this economic transformation is a major challenge [MckinseyReport2022]. For example, to reach net zero by 2050, Shell scenarios suggest that \$3-4 trillion of commercially viable investment in low-carbon energy is required each year up from current \$1.7 trillion a year investment [Shell_energy_2024]. While these are large numbers, the difference from a reasonable baseline energy system investment profile is much smaller.
@@
Generating electricity is a growing contributor to climate change, primarily because coal, oil, and gas are and will still be widely used compared to cleaner energy sources [Shell_Sce_2025]. While electricity demand has been steady over the years, it has recently surged and is projected to continue increasing its share until 2050. This growth is driven by the expanding digital economy and the rise of artificial intelligence (AI). Projections indicate a substantial surge in power consumption from data centers, which are essential for AI infrastructure. Some forecasts suggest that electricity demand from this sector could double by 2026 [iea2024]. The increasing complexity and computational intensity of AI models, especially generative AI, require significantly more energy per query compared to traditional internet services. For instance, a single ChatGPT query consumes approximately ten times the electricity of a typical Google search [deVries2023]. Although the numbers these estimates are based have already been outdated, and the AI assistants can generate more information than just word tokens, it is extensively quoted and generally accepted that queries to AI models consume more energy than simple Google searches. This rising demand puts considerable pressure on existing electricity infrastructure, necessitating a re-evaluation of energy production and consumption patterns [duke2025; epri_2024].
@@
Major tech companies and cloud providers like Microsoft and Google have reported that their emissions have increased recently [googlereport2024; microsoftreport2024]. This trend is expected to continue due to the rush towards building more data centers in the next few years to support the rising AI demand and the ambition to become dominant players in the AI. Several reports, including the IEA report on AI and power demand trends, indicate that GHG emissions will inadvertently rise in the near future as the alternative energy sources are expected to take more time to become significant contributors [Shell_Sce_2025].
@@
However, AI also presents a transformative opportunity for accelerating the transition to a net-zero future. AI’s unique ability to process vast amounts of complex data on energy systems, emissions, and climate impact allows for more informed and data-driven approaches to decarbonization. AI can optimize energy grids, improve the efficiency of renewable energy sources through better forecasting, and support the development of more sustainable industrial processes and materials. Additionally, AI can enhance energy efficiency, availability and reduce waste across various sectors, from transportation and buildings to industrial manufacturing, significantly contributing to emissions reduction. Some reports suggest that AI alone could potentially reduce global GHG emissions by 4% to 16% by 2030 [Shell_Sce_2025; bcg_climate_2022], and 1.4 gigatonnes by 2035 in the widespread adoption case [iea2025].
@@
The advancements in AI for science (see, e.g., [chandra2024fourier]) demonstrate that AI can significantly accelerate complex challenges such as weather modeling and prediction, climate modeling, and CCS storage modeling technologies, potentially by up to 100,000 times. At Shell, we have conducted projects investigating the role of AI in CCS within both traditional aquifers and land use. Our findings indicate that AI can greatly expedite the process of identifying suitable sites and accelerating project execution by several orders of magnitude. We anticipate that in the next 5-10 years, these areas will present significant potential and opportunities to offset the carbon emissions [Pawar2025AcceleratedCS].
@@ 2 Methodology
By employing Monte Carlo simulation with these probabilistically defined inputs, we can generate a range of potential emission scenarios for both the near and long term. This approach allows for a more nuanced understanding of the uncertainties involved and enables sensitivity analysis based on the factors with the most significant influence on future data center and AI-related emissions. The ultimate goal is to provide robust, probabilistic insights that can inform strategies for mitigating the environmental impact of this rapidly growing sector.
@@ 3 The AI Boom and its Energy Appetite
Figure 1 shows the number of foundational models with more than 10 billion parameters built between 2021 and 2024 [data courtesy ecosystemgraph]. It also illustrates the increase in revenues from high-density GPUs from 2022 and the projection for 2028 (Source: electroiq). As can be seen, more investments are being made towards building GPU servers to support the development of large AI models. Although NVIDIA GPUs dominate the market for chips and low-level software libraries, there are no barriers for new entrants to join the arena of building new foundational models. Due to supply and demand mismatches and other approval process-related issues, commissioning grid infrastructure has temporarily slowed the pace of data center growth. However, the overall trend remains one of rapid expansion.
Figure 1: The figure illustrates the growth of foundational models between 2021 and 2024, and the revenue growth of high-density GPU markets. Data courtesy from ecosystemgraph
@@
However, this progress comes with significant consequences: a rapidly increasing demand for computational resources required to support the tasks in the life cycle of AI models, which involves both training (a one-time or occasional effort) and inference (repeatedly executed when the model is deployed for consumption). Although the work done (measured in FLOPs) per dollar and per watt has increased, meaning the power consumption per FLOP has decreased (see Figure 2 (left)), the number of FLOPS required to train AI models has continued to rise (see Figure 2 (right)). Consequently, the total power consumption for training AI models keeps increasing. This translates directly into escalating energy consumption [goldman2025]. For instance, it was reported in [PattersonGoogle2021] that training the large language model (LLM) powering ChatGPT-3 consumed almost 1300 megawatt hours of energy, equivalent to the annual energy usage of 130 American homes. Furthermore, an analysis by OpenAI suggests that the amount of power needed to train AI models has grown exponentially since 2012, doubling roughly every 3.4 months as the models become more sophisticated [openaiblog].
@@
Computational Demands of Training and Inference: AI models, especially deep learning models and LLMs, require immense computational power for both training and inference [deVries2023].
@@
  • Inference is the deployment and utilization of the trained AI models in real-world applications, where the model interprets new data and generates outcomes or predictions. Despite innovations in model efficiency, the inference phase accounts for the majority (60%) of the energy footprint [epri_2024] (see Table 1 for electricity consumption of various tasks that AI chatbots are typically asked to execute).
@@ 4 Data Centers: The Epicenter of AI’s Energy Demand
Data centers are the backbone of the AI ecosystem, serving as the primary energy consumers that power the computational workloads required for training and deploying AI models. These facilities house vast arrays of servers, storage devices, and networking equipment, all of which demand significant amounts of electricity to operate. The rapid expansion of AI, particularly deep learning and large language models, has led to an unprecedented surge in data center growth, intensifying concerns about their environmental impact [epri_2024].
The data center industry is experiencing rapid growth worldwide, driven by the increasing demand for cloud computing, digital services, cryptocurrencies, and AI applications. The number of data centers is increasing, and their sizes are also expanding. As of March 2024, there were approximately 9580 data centers globally, with more than half of them located in the United States [datacentermap]. Between 2015 and 2024, the electricity consumption at data centers increased from under 200 TWh to over 400 TWh [iea2025]. Data centers are projected to consume between 4.6% and 9.1% of U.S. electricity generation annually by 2030, up from an estimated 4% today. Some forecasts even predicted that data centers could draw up to 21% of the world’s electricity supply by 2030 [Jones2018]. In the US, fifteen states account for 80 percent of the national data center load, with data centers comprising a quarter of Virginia’s electric load in 2023. New data centers are being built with capacities ranging from 100 to 1000 megawatts, equivalent to the load from 80,000 to 800,000 homes.
In [epri_2024], the authors have developed four growth scenarios for US data center electricity load from 2023 to 2030: low growth (3.7 percent annual), moderate growth (5 percent annual), high growth (10 percent annual), and higher growth (15 percent annual). In the higher growth scenario, data center loads are projected to increase by a staggering 166% from 152 TWh/year in 2023 to 403 TWh/year in 2030, with the majority of this demand attributed to AI. Such projections highlight the significant energy demands of AI and data centers and the need for innovative energy management and infrastructure planning to accommodate this growth.
@@
Despite the growing emphasis on renewable energy, many regions still rely heavily on fossil fuels, especially, natural gas, to power their electricity grids [mittechreview2025]. This may mean that a significant portion of data centers’ energy consumption is directly contributing to GHG emissions. In regions with limited access to clean energy sources, the increasing electricity demands of data centers can exacerbate the reliance on fossil fuels, hindering efforts to decarbonize the energy sector. For instance, in [mittechreview2025], the author quoted that natural gas is the only affordable choice given the need to meet the 24-7 electricity demand from the huge data center.
@@
As AI has become increasingly pervasive, it is difficult to estimate how much electricity consumption can be attributed solely to AI-related workloads. Some studies using proxies estimate that AI accounts for 10% to 20% of total data center electricity consumption [see iea2025 for example], but this percentage is rapidly increasing. AI workloads, including computation, storage, and networking, have specific energy needs within data centers [epri_2024; iea2024].
@@ 5 The Energy Consumption Scenarios and Near-Term Outlook for Data Centers and AI
Shell’s 2025 Energy Security Scenarios explore AI’s transformative role across different pathways, each highlighting distinct impact on energy systems and societal change [Shell_Sce_2025]. In the Surge scenario, AI is a key driver of economic growth and societal transformation. AI optimizes manufacturing processes, enhances productivity across various sectors, and leads to increased economic activity. This surge in economic activity results in higher energy demand, which is managed efficiently by AI-powered smart grids. Additionally, AI accelerates the transition to low-carbon energy sources by optimizing the integration of renewable energy and enhancing energy storage solutions. The Archipelagos scenario presents a different picture, where AI development is hindered by global concerns about resource security, border control, and trade restrictions. Geopolitical tensions limit the sharing of AI technologies and data, slowing progress. As a result, technological advancements are slower, impacting the transformation of energy systems. In the Horizon scenario, AI plays a crucial role in achieving net-zero emissions by 2050. AI-driven innovations in energy efficiency and carbon capture technologies are pivotal in reaching this goal. AI also helps limit global warming to 1.5°C by optimizing energy consumption and reducing GHG emissions (Figure 4).
Figure 4: Electricity generation and associated CO2 emissions between 2020-2060 under the three Shell Energy Security scenarios. Of those, only Horizon is reaching Net Zero by the 2050s. [Courtesy of Shell_Sce_2025].
@@
Building on the insights from Shell’s Energy Securities scenarios, the Energy and AI 2025 report by the International Energy Agency (IEA) [iea2025] outlines four sensitivity cases to explore different scenarios for AI uptake and its impact on the energy sector (Figure 5a). The Lift-Off Case assumes higher rates of AI adoption and proactive measures to reduce energy sector bottlenecks, envisioning rapid expansion of AI technologies and infrastructure. This scenario highlights the transformative potential of AI when supported by robust energy policies and investments, leading to significant advancements in AI capabilities and widespread adoption. The Headwinds Case incorporates various bottlenecks, including macroeconomic challenges and delays in building out necessary energy infrastructure, representing a more cautious and constrained growth scenario. This case underscores the importance of addressing external factors that can hinder AI development, such as economic instability and infrastructure limitations, which can slow progress and reduce the overall impact of AI.
@@
Figure 5: a) Four scenarios for electricity demand by data centers globally between 2020-2035 (reproduced from iea2025); b) Plausible scenarios for electricity usage by AI applications between 2020 and 2035. (reproduced from paccouandwijnhoven2024).
@@
In a recent study [paccouandwijnhoven2024], the authors presented a comprehensive outlook using a system dynamics approach to model four scenarios for AI electricity consumption: Sustainable AI, Limits To Growth, Abundance Without Boundaries, and Energy Crisis. These scenarios provide insights into potential future trajectories of AI electricity use, influenced by current AI infrastructure decisions. While all scenarios show initial growth, their paths diverge after 2030 due to various internal and external factors (Figure 5b). The Sustainable AI scenario prioritizes efficiency and sustainability, with energy consumption increasing from 100 TWh in 2025 to 785 TWh in 2035. This scenario balances technological advancement with environmental stewardship. In contrast, the Limits To Growth scenario shows minimal increase in consumption, from 510 TWh in 2030 to 570 TWh in 2035, indicating stunted economic expansion due to constraints like power availability and data scarcity. The Abundance With- out Boundaries scenario projects unchecked growth, with consumption reaching 1,370 TWh by 2035. This scenario highlights risks such as power centralization and inequitable AI access. Meanwhile, the Energy Crisis scenario underscores the risks of mismatched energy demand and infrastructure, with consumption peaking at 670 TWh before dropping to 190 TWh by 2035, indicating potential energy crisis.
@@
Specific insights from these scenarios include the dominance of generative AI in the Sustainable AI scenario, focusing on less energy-intensive models. The Limits To Growth scenario highlights constraints in AI development due to power, chip manufacturing, and data scarcity. The Abundance Without Boundaries scenario warns of risks associated with unchecked AI growth, including unsustainable infrastructure costs, global inequality, and increased e-waste. The Energy Crisis scenario emphasizes the likelihood of underestimated future electricity needs due to insufficient grid planning and inaccurate demand projections. The authors note that projections for electricity consumption in TWh face challenges due to the nascent nature of the discipline, making direct comparisons difficult. However, some convergence is observed: the AI Abundance Without Boundaries scenario projects 880 TWh usage by 2030, aligning with estimates by [pateletal2024]. Another projection by [Kindig2024] of 652 TWh by 2030 aligns with the Sustainable AI scenario, suggesting a balanced growth trajectory.
@@
To provide a comprehensive understanding of the potential futures for AI development and energy consumption, we have matched the Shell scenarios with the IEA sensitivity cases and the AI scenarios from paccouandwijnhoven2024. This alignment is based on the shared themes and underlying assumptions of each scenario. The Shell scenarios offer a broad perspective on global economic and technological trends, while the IEA sensitivity cases focus on the specific impacts of AI on the energy sector. The AI scenarios from paccouandwijnhoven2024 provide detailed insights into the potential trajectories of AI electricity consumption. By matching these scenarios, we can explore how different factors such as geopolitical tensions, technological advancements, and sustainability efforts might influence the future of AI and energy. This integrated approach allows us to identify common challenges and opportunities, offering a more nuanced and holistic view of the interplay between AI development and energy consumption:
@@
    Surge - Lift-Off Case - Abundance Without Boundaries: The Surge scenario envisions rapid AI adoption and technological advancements, driving significant economic growth. Similarly, the Lift-Off Case from the IEA assumes high AI adoption rates and proactive measures to reduce energy bottlenecks, highlighting AI’s transformative potential. The Abundance Without Boundaries scenario by paccouandwijnhoven2024 projects unrestricted AI growth, emphasizing its power to drive economic and technological progress without major constraints. These scenarios share a vision of rapid and expansive AI growth fueling economic and technological advancements. Under these scenarios, the energy demand by data centers continues to grow past 2035 exceeding 1700 TWh. AI accounts for more than three quarters of the electricity usage (Figure 6a).
@@
    Archipelagos - Headwinds Case - Limits To Growth: The Archipelagos scenario depicts technology development hindered by global concerns about resource, border, and trade security, leading to fragmented efforts and slower progress. Similarly, the Headwinds Case from the IEA includes various bottlenecks, such as macroeconomic challenges and infrastructure delays, representing cautious growth with significant obstacles. The Limits To Growth scenario by paccouandwijnhoven2024 highlights constraints on AI development due to resource limitations, emphasizing the need for careful management of AI’s energy demands. These scenarios share a focus on limitations and challenges that hinder AI and energy sector progress. Under these scenarios, energy demand by data centers plateaus by 2030 at around 700 TWh. AI accounts for approximately 80% of the electricity usage (Figure 6b).